OpenStudy (cookiemonster18):

1.Tell whether the sequence is arithmetic. If it is, what is the common difference? -19,-11,-3,5... A.yes;5 B.yes;6 C.yes;8 D.no 2.What is the slope of that passes through the pair of points? (-6,8),(2,3) A.-3/8 B.-1/8 C.-7/8 D.-5/8 3.Write an equation in point-slope form for the line through the given point with the given slope. (-3,5);m=2/3 A.y+3=2/3(x-5) B.y+5=(2/3)(x+3) C.y-5=2/3(x-3) D.y-5=2/3(x+3) 4.Write y=-2/5x-3 in standard form using integers. A.2x+5y=-15 B.2x+y=-3 C.2x+5y=-3 D.x+y=-15

OpenStudy (cookiemonster18):

5.Write the equation of a line that is perpendicular to the given line and that passes through the given point. y-2=-1/4(x+3);(-3,5) A.y=-3x-5 B.y=1/4x-6 C.y=4x+17 D.y=4x+5 6.How many solutions doe the system of equations have? 3x+15y=9 y=-1/5x+9/15 A.one B.two C.infinitely many D.none I think #1.D #2.C #3.A #4.B #5.A #6.A (Pleas correct me if im wrong)

OpenStudy (anonymous):

One question at a time, please.

OpenStudy (anonymous):

Question 1: -19,-11,-3,5... \[-19 + (8) = -11\] \[-11 + (8) = -3\] \[-3 + (8) = 5\] As you may have noticed, this is an arithmetic with a consistent increment of 8.